OpenStudy (anonymous):

Theoretical example system of DEs with only 1 eigenvector and eigenvalues mu_g(lambda)

OpenStudy (anonymous):

\[x ^{'}(t)=\left[\begin{matrix}1 & -1 \\ 1 & 3\end{matrix}\right]x(t)\] eigenvalue = 2 eigenvector = (1,-1) \[\rightarrow x _{1}(t)=\left(\begin{matrix}1 \\ -1\end{matrix}\right)e ^{2t}\] Problem: only one apparent solution where we expect 2. Simple calculus method: replace e^2t with te^2t and find solution to \[x _{2}(t)=\eta te ^{2t}\] \[x^{'} _{2}(t)=2\eta te ^{2t} + \eta e ^{2t}\] \[\rightarrow 2\eta te ^{2t} + \eta e ^{2t} - \left[\begin{matrix}1 & -1 \\1 & 3\end{matrix}\right]\eta te ^{2t}=0\] which could only work if eta = 0 Since the last equation holds both e^2t as well as te^2t we could combine them and say that \[x _{2}(t)=\eta te ^{2t} + \zeta e ^{2t}\] after taking the derivative that would form the following equation \[2\eta te ^{2t}+ (\eta + 2\zeta) e ^{2t}=\left[\begin{matrix}1 & -1 \\ 1 & 3\end{matrix}\right](\eta te ^{2t} + \zeta e ^{2t})\] I follow this far. But then my course says "If we call the matrix P and take the coefficients in e^2t en te^2t together then we find \[(P-21_{2x2})\eta=0\] \[(P-21_{2x2})\zeta=\eta\] The 1 of 21 in these equations are boldened. Can someone explain me where these equations come from?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

It's the last two equations with P that I don't get.

OpenStudy (anonymous):

I have time :-)

OpenStudy (anonymous):

The course continues by saying that the first equation tells us that eta=(1,-1)^T, but there is something strange going on with the second: the determinant of the matrix equals zero, which can make us think there is no solution to the second equation (simply inverting to get a solution is not an option). For some eta's one could solve the equation: "the vectors eta that are perpendicular (Euclidian inprouct) to various freely chosen y's for the system \[(P ^{\dagger} - 21_{2x2})y=0\] With P^dagger the Hermitic addition. So, this all serves to claim that zeta does exist. They can be found with the system \[-\zeta _{1} -\zeta _{2}=+1\] \[+\zeta _{1} +\zeta _{2}=-1\] The idea behind the above condition about eta is that to say that eta (element of R^2) must be perpendicular to vectors y in ker M^dagger is like saying that \[R ^{2}= \ker M circledcross \ker M^\dagger\] with \[M = (P - 21_{2x2}\]

OpenStudy (anonymous):

Yup the second eigenvector in the course ends up being (0,-1). But I just don't get the whole stuff about (P-21_2x2). I don't really understand its meaning and where my prof gets it from.

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Why did you mention me? What is the question?

OpenStudy (anonymous):

Are you sure it isn't \[ P - 2I_{2\times 2} = P -\begin{bmatrix} 2&0\\ 0&2 \end{bmatrix} \]

OpenStudy (anonymous):

I need someone to explain to me what the (P-...) expressions and equations mean and where they come from.

OpenStudy (anonymous):

It's a boldened 1, not an I

OpenStudy (anonymous):

Well you have to look it up in your book then, because it's not standard notation.

OpenStudy (anonymous):

I've looked all over the linear algebra course, dagger and ker I find, but not that notation. :(

OpenStudy (anonymous):

Different people will say it means different things.

OpenStudy (anonymous):

wikipedia will say it is identity matrix, others wills say is it vector of ones.

OpenStudy (anonymous):

I've put the question to my professor by email, to make sure it's capital I for Identity matrix or not.

OpenStudy (anonymous):

Thanks at least for suggesting an idea about it, @wio

OpenStudy (anonymous):

The purpose is to get a 0 determinate, usually

OpenStudy (anonymous):

Which will ensure the null space is a vector, your Eigen vector.

OpenStudy (anonymous):

Yeah, I'm starting to get it now... He probably used 1 to indicate it's a vector of ones on the diagonal, but it's less clean that using identity notation.

OpenStudy (anonymous):

the -2I_2x2 is the eigenvalue in matrix notation form.